OpenStudy (anonymous):

x + 2y = 10 3x + 4y = 8 Which point is the solution to the system of equations? A)(7, 1.5) B)(11,-12) C)(-11, -12) D)(-12, 11)

OpenStudy (anonymous):

@Spectrum

OpenStudy (anonymous):

Solve for x in the first equation and then plug in that solution into the x in the second equation. Then just solve for the y-value. Then plug that y-value into any of the two equations and solve for x.

OpenStudy (anonymous):

huuuh ?

OpenStudy (anonymous):

Solve for x in x+2y=10

OpenStudy (anonymous):

how ? by subtracting 2Y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

X is understood to be 1

OpenStudy (anonymous):

@Twis7ed

OpenStudy (anonymous):

What? Just subtract 2y from both sides so you end up with x = 10 - 2y. Do you understand that?

OpenStudy (anonymous):

right i got that well then you have to divide by the number infront of X and there isnt one so itnt it understood to be 1

OpenStudy (anonymous):

yes, and you have the value for x which is (10-2y). Use (10-2y) in the second equation. that is what he is saying.

OpenStudy (anonymous):

so it would be like 3(10-2y)+4y=8

OpenStudy (anonymous):

@franciscanmonk

OpenStudy (anonymous):

omg i need someone whho makes sence like @some_someone or @bahrom7893 or @Spectrum

OpenStudy (bahrom7893):

Do I make sense?

OpenStudy (anonymous):

yes , you did in my last question you answered

OpenStudy (anonymous):

'so it would be like 3(10-2y)+4y=8' <--- This one is good. Solve this get the value for y solve for x in either the first or second equation.

OpenStudy (bahrom7893):

x+2y=10 Subtract 2y from both sides: x+2y -2y = 10 -2y

OpenStudy (anonymous):

got that baharom ! so i multiply the 3 by 10-2y ?

OpenStudy (bahrom7893):

So now you have: x=10-2y Since you know that 3x + 4y = 8 and x=10-2y, you know that: 3*(10-2y) + 4y = 8

OpenStudy (anonymous):

alright so that equals 30-6y+4y=8 then i combine like terms which are 6y and 4y which equal 10y ?

OpenStudy (bahrom7893):

yea so 30-6y+4y=8 30-2y=8 -2y=8-30 -2y=-22 y=11

OpenStudy (anonymous):

X=12 right ?

OpenStudy (anonymous):

answer is 11,-12

OpenStudy (bahrom7893):

x=10-2y = 10-2*11 = 10-22 = -12

OpenStudy (anonymous):

its D

OpenStudy (bahrom7893):

no

OpenStudy (bahrom7893):

In coordinate form you write (x,y) so it's B

OpenStudy (anonymous):

i clicked b , it said i was wrong